OpenStudy (anonymous):

whats the answer to 10 2/25 - 4 1/20 =

OpenStudy (anonymous):

no its a question

OpenStudy (anonymous):

Convert the fractions into improper fractions, find a common denominator, subtract and then simplify.

OpenStudy (anonymous):

okay

OpenStudy (radar):

10-4=6 8/100-5/100=3/100 answer is 6 3/100